# Classical harmonic analysis viewed through the prism of noncommutative geometry

### v3

@@ 8 Examples of triples and Banach Fredholm modules
of [BEO13, Corollary 1.2] , which gives a bounded embedding $J\colon\mathrm{L}^{p}_{\alpha}(\mathbb{T}^{d})\to\mathrm{L}^{q}(\mathbb{T}^{d})$. Note that Sobolev embeddings are also available for compact Riemannian manifolds, see e.g., [ReS24, Proposition 2.2]. Moreover, if $1\leqslant p<q\leqslant 2$ satisfy $\frac{\alpha}{d}>\frac{1}{p}-\frac{1}{q}$ the result [Kon86, Theorem pp. 186-187] provides essentially the estimate
(8.4) $a_{n}(J)\lesssim\frac{1}{n^{\frac{\alpha}{d}}},\quad n\geqslant 1$
on the approximation numbers.
@@
Suppose that $1<p<\infty$. Then $(\mathrm{C}(\mathbb{T}),\mathrm{L}^{p}(\mathbb{T}),\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta})$ is a Banach compact spectral triple which is $q$-summable for any $q>1$.
Proof : Note that it is well known [AGGG+86, pp. 10-11] that the operator $\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}$ generates a strongly continuous group of operators acting on $\mathrm{L}^{p}(\mathbb{T})$, namely the group of translations. By Example 2.7 and Example 2.8, we deduce that the unbounded operator $D\overset{\mathrm{def}}{=}\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}$ is bisectorial with $\omega_{\mathrm{bi}}(D)=0$ and admits a bounded $\mathrm{H}^{\infty}(\Sigma_{\omega}^{\mathrm{bi}})$ functional calculus for any angle $\omega>0$ on the Banach space $\mathrm{L}^{p}(\mathbb{T})$. For any function $f\in\mathrm{C}^{\infty}(\mathbb{T})$, we have $M_{f}\cdot\operatorname{dom}\tfrac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}\subset\operatorname{dom}\tfrac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}$ since $\operatorname{dom}\tfrac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}=\mathrm{W}^{1,p}(\mathbb{T})$ and an elementary computation reveals that
(8.5) $\big[\tfrac{1}{\mathrm{i}}\tfrac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta},f\big]=\tfrac{1}{\mathrm{i}}\big(\tfrac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}M_{f}-M_{f}\tfrac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}\big)=\frac{1}{\mathrm{i}}M_{f^{\prime}}.$
So the commutator $\frac{1}{\mathrm{i}}\tfrac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta},f]$ induces a bounded operator on the Banach space $\mathrm{L}^{p}(\mathbb{T})$. Finally, it is stated without proof in [Kat76, Example 6.31 p. 187] that the operator $\frac{1}{\mathrm{i}}\tfrac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}$ has compact resolvent. Definition 7.4 is satisfied. Actually, below we prove a stronger fact than the stated result in [Kat76, Example 6.31 p. 187].
Suppose that $1<p<2$. Consider some $q\in(1,p)$ and some $\alpha>0$. For any function $f\in\mathrm{L}^{p}(\mathbb{T})$, we have
@@
If $\alpha>\frac{1}{q}-\frac{1}{p}$, using the Sobolev inequality (8.3), we deduce the factorization
(8.6) $(\mathrm{Id}+\Delta)^{-\frac{\alpha}{2}}\colon\mathrm{L}^{p}(\mathbb{T})\xrightarrow{(\mathrm{Id}+\Delta)^{-\frac{\alpha}{2}}}\mathrm{L}^{q}_{\alpha}(\mathbb{T})\xrightarrow{J}\mathrm{L}^{p}(\mathbb{T}).$
For any integer $n\geqslant 1$, we obtain
@@
Consequently, the series $\sum a_{n}((\mathrm{Id}+\Delta)^{-\frac{\alpha}{2}})$ converges if $\alpha>1$. We deduce that the linear operator $(\mathrm{Id}+\Delta)^{-\frac{\alpha}{2}}\colon\mathrm{L}^{p}(\mathbb{T})\to\mathrm{L}^{p}(\mathbb{T})$ belongs to the space $S^{1}_{\mathrm{app}}(\mathrm{L}^{p}(\mathbb{T}))$, defined in (5.7). Thus the triple $(\mathrm{C}(\mathbb{T}),\mathrm{L}^{p}(\mathbb{R}),\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta})$ is $q$-summable for any $q>1$. If $p>2$, we reason by duality using (5.4). The case $p=2$ is well known.
Remark 8.2
Suppose that $1<p<\infty$. In this remark, we sketch a proof showing that the previous result is optimal. When $p=2$, the statement is well known. The Banach compact spectral triple $(\mathrm{C}(\mathbb{T}),\mathrm{L}^{p}(\mathbb{T}),\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta})$ is not $1$-summable. Indeed, by [AmG16, Remark B.7 p. 127], the spectrum of the Dirac operator $D\overset{\mathrm{def}}{=}\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}$ is independent of $p$. In the case $p=2$, it is known that $\operatorname{Sp}D=\pi\mathbb{Z}$ and that each eigenvalue is simple, see e.g., [Var10, Exercise 8.1] or [Gin09, Notes 2.1.2 p. 31]. Using Weyl’s inequality (5.10) and a ¡¡bisectorial version¿¿ (left to the reader) of the spectral theorem of [Haa06, p. 56] [Haa05], we see that for any $z\not\in\operatorname{Sp}D$ we have
$\infty=\sum_{n\in\mathbb{Z}}\frac{1}{|z-\pi n|}=\sum_{n\in\mathbb{Z}}\lambda_{n}(R_{z}(D))=\left\|\big(\lambda_{n}(R_{z}(D))\big)\right\|_{\ell^{1}}$
@@
where we use the resolvent operator $R_{z}(D)\overset{\mathrm{def}}{=}(z-D)^{-1}$. Consequently the operator $R_{z}(D)$ does not belong to the space $S^{1}_{\mathrm{app}}(\mathrm{L}^{p}(\mathbb{T}))$. We conclude that the triple is not $1$-summable according to Definition 7.3.
@@
(8.10) $\big\langle[f],(\mathrm{L}^{p}(\mathbb{T}),\pi,\mathrm{i}\mathcal{H}_{\mathrm{per}})\big\rangle_{\mathrm{K}_{1}(\mathrm{C}(\mathbb{T})),\mathrm{K}^{1}(\mathrm{C}(\mathbb{T}),\mathscr{L}^{p})}=-\operatorname{wind}f,$
@@
Proof : The operator $|D|$ is the unbounded Fourier multiplier defined by the symbol $(|n|)_{n\in\mathbb{Z}}$ and $D$ by $(n)_{n\in\mathbb{Z}}$. The first assertion is then a consequence of (8.8). Hence, for any function $g\in\mathrm{C}^{\infty}(\mathbb{T})$, we have
@@
We now estimate the Hölder differences. For the first one, by the fundamental theorem of calculus,
@@
Suppose that $1<p<\infty$. Note that by [EnN00, Proposition 4.10 p. 31] any multiplication operator $M_{f}\colon\mathrm{L}^{p}(\mathbb{R})\to\mathrm{L}^{p}(\mathbb{R})$, $g\mapsto fg$ by a measurable function $f\colon\mathbb{T}\to\mathbb{C}$ is bounded if and only if $f$ is essentially bounded. In this case, we have $\left\|M_{f}\right\|_{\mathrm{L}^{p}(\mathbb{R})\to\mathrm{L}^{p}(\mathbb{R})}=\mathrm{L}^{\infty}(\mathbb{R})$. Consequently, we can consider the isometric homomorphism $\pi\colon\mathrm{L}^{\infty}(\mathbb{R})\to\mathrm{B}(\mathrm{L}^{p}(\mathbb{R}))$, $f\mapsto M_{f}$. It is well known [Ren04, Corollary 14 p. 92] that $(\mathrm{C}_{0}(\mathbb{R}),\mathrm{L}^{2}(\mathbb{R}),\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x})$ is a locally compact spectral triple, where the used homomorphism is the restriction of $\pi$ on the algebra $\mathrm{C}_{0}(\mathbb{R})$. First, we prove a weaker $\mathrm{L}^{p}$-variant of this classical fact. Here, we consider the closure of the unbounded operator $\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}\colon\mathrm{C}^{\infty}_{0}(\mathbb{R})\subset\mathrm{L}^{p}(\mathbb{R})\to\mathrm{L}^{p}(\mathbb{R})$. We denote again $\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}$ its closure whose the domain $\operatorname{dom}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}$ is the Sobolev space
$\mathrm{W}^{1,p}(\mathbb{R})\overset{\mathrm{def}}{=}\{f\in\mathrm{L}^{p}(\mathbb{R}):f\text{ is absolutely continuous},f^{\prime}\in\mathrm{L}^{p}(\mathbb{R})\}.$

### v4

@@ 8 Examples of triples and Banach Fredholm modules
of [BEO13, Corollary 1.2], which gives a bounded embedding $J\colon\mathrm{W}^{\alpha,p}(\mathbb{T}^{d})\to\mathrm{L}^{q}(\mathbb{T}^{d})$. Note that Sobolev embeddings are also available for compact Riemannian manifolds, see e.g., [ReS24, Proposition 2.2]. We need the following estimates on the approximation numbers.
Proposition 8.1
Suppose that $1<p\leqslant q\leqslant 2$. Consider some $\alpha>0$ with $\frac{\alpha}{d}>\frac{1}{p}-\frac{1}{q}$. If $J\colon\mathrm{W}^{\alpha,p}(\mathbb{T}^{d})\to\mathrm{L}^{q}(\mathbb{T}^{d})$ is the canonical Sobolev embedding then
(8.2) $a_{n}(J)\lesssim\frac{1}{n^{\frac{\alpha}{d}}},\qquad n\geqslant 1.$
Proof : Consider the restriction map $R_{p}\colon\mathrm{L}^{p}_{\alpha}(\mathbb{T}^{d})\to\mathrm{W}^{\alpha,p}((0,2\pi)^{d})$, $f\mapsto f|_{(0,2\pi)^{d}}$ and the periodization $\mathcal{P}_{q}\colon\mathrm{L}^{q}((0,2\pi)^{d})\to\mathrm{L}^{q}(\mathbb{T}^{d})$. Both maps are bounded. The range $X_{p}\overset{\mathrm{def}}{=}\operatorname{Ran}R_{p}$ is a closed subspace of $\mathrm{W}^{\alpha,p}((0,2\pi)^{d})$. Consider the inclusion $j\colon X_{p}\hookrightarrow\mathrm{L}^{q}((0,2\pi)^{d})$. By construction, the Sobolev embedding on $\mathbb{T}^{d}$ factorizes as
$J=\mathcal{P}_{q}\circ j\circ R_{p}.$
The ideal (composition) property of approximation numbers yields
[…]
Since $X_{p}$ is a closed subspace of $\mathrm{W}^{\alpha,p}((0,2\pi)^{d})$ and restriction of the domain can only decrease approximation numbers, we have
$a_{n}(j)\leqslant a_{n}(J_{0}),$
where $J_{0}\colon\mathrm{W}^{\alpha,p}((0,2\pi)^{d})\hookrightarrow\mathrm{L}^{q}((0,2\pi)^{d})$ denotes the Sobolev embedding. By [Kon86, Theorem pp. 186–187], we have $a_{n}(J_{0})\lesssim\frac{1}{n^{\frac{\alpha}{d}}}$. We conclude that (8.2) is true.
We also need the next result.
Proposition 8.2
Let $(T_{t})_{t\in\mathbb{R}}$ be a strongly continuous group of isometries on a Banach space $X$ with generator $\mathrm{i}D$, i.e., $T_{t}=\mathrm{e}^{\mathrm{i}tD}$ for any $t\in\mathbb{R}$. Consider a bounded operator $R\colon X\to X$ satisfying $R(\operatorname{dom}D)\subset\operatorname{dom}D$ such that $[D,R]$ induces a bounded operator on $X$. Then the function $F\colon\mathbb{R}\to\mathrm{B}(X)$ defined by
$F(t)=T_{t}RT_{-t},\quad t\in\mathbb{R}.$
is strongly $\mathrm{C}^{1}$ and
(8.3) $F^{\prime}(t)=\mathrm{i}T_{t}[D,R]T_{-t}\qquad\text{(strongly on $X$)}$
and
(8.4) $\frac{T_{t}RT_{-t}-R}{t}=\int_{0}^{1}T_{st}(\mathrm{i}[D,R])T_{-st}\mathop{}\mathopen{}\mathrm{d}s.$
Proof : Fix $x\in\operatorname{dom}D$. Using the generator identities [EnN00, Lemma 1.3 p. 50]
$\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}t}T_{t}x=\mathrm{i}DT_{t}x,\qquad\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}t}T_{-t}x=-\mathrm{i}T_{-t}Dx\quad(\text{derivatives in $X$}),$
and the product rule in the strong sense, we get
(8.5) $\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}t}\big(T_{t}RT_{-t}x\big)=\mathrm{i}DT_{t}RT_{-t}x+T_{t}R(-\mathrm{i})T_{-t}Dx=\mathrm{i}T_{t}\big(DR-RD\big)T_{-t}x=\mathrm{i}T_{t}[D,R]T_{-t}x.$
By density of $\operatorname{dom}D$ and boundedness of $[D,R]$, this extends to all $x\in X$, proving (8.3).
Now fix $t\in\mathbb{R}$ and $x\in X$. The map $u\mapsto T_{u}[D,R]T_{-u}x$ is continuous (as a vector-valued function) and bounded by $\left\|[D,R]\right\|\left\|x\right\|$, hence Bochner integrable on $[0,t]$. The Banach-valued fundamental theorem of calculus yields
[…]
Dividing by $t\neq 0$ and using the change variables $u=st$, we obtain
$\frac{T_{t}RT_{-t}-R}{t}x=\int_{0}^{1}T_{st}(\mathrm{i}[D,R])T_{-st}x\mathop{}\mathopen{}\mathrm{d}s.$
Since this holds for every $x\in X$, we have in the strong operator topology the operator identity
$\frac{T_{t}RT_{-t}-R}{t}=\int_{0}^{1}T_{st}(\mathrm{i}[D,R])T_{-st}\mathop{}\mathopen{}\mathrm{d}s.$
Now, we can determine the Lipschitz algebra of the operator $\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}$. This result is new even for $p=2$.
Proposition 8.3
Suppose that $1<p<\infty$. Consider the operator $D\overset{\mathrm{def}}{=}\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}$ with domain $\operatorname{dom}D=\mathrm{W}^{1,p}(\mathbb{T})$. Then
(8.6) $\mathrm{W}^{1,\infty}(\mathbb{T})=\operatorname{\mathrm{Lip}}_{D}(\mathrm{L}^{\infty}(\mathbb{T})).$
Moreover, for any function $f\in\mathrm{W}^{1,\infty}(\mathbb{T})$ we have
(8.7) $D,M_{f}]=\frac{1}{\mathrm{i}}M_{f^{\prime}}\quad\text{and}\quad\left\|[D,M_{f}]\right\|_{\mathrm{L}^{p}(\mathbb{T})\to\mathrm{L}^{p}(\mathbb{T})}=\left\|f^{\prime}\right\|_{\mathrm{L}^{\infty}(\mathbb{T})}.$
Proof : If $f\in\mathrm{W}^{1,\infty}(\mathbb{T})$ then $M_{f}(\mathrm{W}^{1,p}(\mathbb{T}))\subset\mathrm{W}^{1,p}(\mathbb{T})$. Moreover, for any function $u\in\mathrm{C}^{\infty}(\mathbb{T})$ we have
$D,M_{f}]u=(DM_{f}-M_{f}D)u=\tfrac{1}{\mathrm{i}}\big((fu)^{\prime}-fu^{\prime}\big)=\tfrac{1}{\mathrm{i}}f^{\prime}u.$
Hence $D,M_{f}]=\tfrac{1}{\mathrm{i}}M_{f^{\prime}}$ extends boundedly on $\mathrm{L}^{p}(\mathbb{T})$ with $\left\|[D,M_{f}]\right\|_{\mathrm{L}^{p}(\mathbb{T})\to\mathrm{L}^{p}(\mathbb{T})}=\left\|f^{\prime}\right\|_{\mathrm{L}^{\infty}(\mathbb{T})}$. In particular, we obtained $\mathrm{W}^{1,\infty}(\mathbb{T})\subset\operatorname{\mathrm{Lip}}_{D}(\mathrm{L}^{\infty}(\mathbb{T}))$.
Note that it is well known [AGGG+86, pp. 10-11] that the operator $\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}$ generates a strongly continuous group $(T_{t})_{t\in\mathbb{R}}$ of operators acting on $\mathrm{L}^{p}(\mathbb{T})$, namely the group of translations. For any $t\in\mathbb{R}$ we have $T_{t}=\mathrm{e}^{\mathrm{i}tD}$ and
(8.8) $T_{t}M_{f}T_{-t}=M_{f(\cdot+t)},\qquad t\in\mathbb{R}.$
Let $f\in\operatorname{\mathrm{Lip}}_{D}(\mathrm{L}^{\infty}(\mathbb{T}))$. Since $D,M_{f}]\in\mathrm{B}(\mathrm{L}^{p}(\mathbb{T}))$ by Proposition 8.2 we deduce that
[…]
If $|t|$ is small, we infer that
[…]
$\leqslant\int_{0}^{1}\left\|T_{st}(\mathrm{i}[D,M_{f}])T_{-st}\right\|_{\mathrm{B}(\mathrm{L}^{p}(\mathbb{T}))}\mathop{}\mathopen{}\mathrm{d}s\leqslant\left\|[D,M_{f}]\right\|_{\mathrm{B}(\mathrm{L}^{p}(\mathbb{T}))}.$
By Banach–Alaoglu theorem, there exist $t_{\alpha}\to 0$ and $g\in\mathrm{L}^{\infty}(\mathbb{T})$ such that $\Delta_{t_{\alpha}}f\overset{}{\longrightarrow}g$ in the weak* topology of the dual space $\mathrm{L}^{\infty}(\mathbb{T})$, with $\left\|g\right\|_{\mathrm{L}^{\infty}(\mathbb{T})}\leqslant\left\|[D,M_{f}]\right\|_{\mathrm{B}(\mathrm{L}^{p}(\mathbb{T}))}$. For any function $\varphi\in\mathrm{C}^{\infty}(\mathbb{T})$, we have with a change of variables
$\int_{\mathbb{T}}(\Delta_{t_{\alpha}}f)\varphi=\int_{\mathbb{T}}f\frac{\varphi(\cdot-t_{\alpha})-\varphi}{t_{\alpha}}\xrightarrow[\alpha]{}-\int_{\mathbb{T}}f\,\varphi^{\prime}$
since $\frac{\varphi(\cdot-t)-\varphi}{t}\to-\varphi^{\prime}$ in $\mathrm{L}^{1}(\mathbb{T})$ as $t\to 0$ (since $\mathrm{i}D$ is the generator of $(T_{t})_{t\in\mathbb{R}}$ on $\mathrm{L}^{1}(\mathbb{T})$). Passing to the weak* limit gives
$\int_{\mathbb{R}}g\varphi=-\int_{\mathbb{R}}f\varphi^{\prime},\quad\varphi\in\mathrm{C}^{\infty}(\mathbb{T}),$
i.e., $g$ is the weak derivative of $f$. Hence $f$ belongs to $\mathrm{W}^{1,\infty}(\mathbb{T})$.
@@
Suppose that $1<p<\infty$. Then $(\mathrm{C}(\mathbb{T}),\mathrm{L}^{p}(\mathbb{T}),\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta})$ is a Banach compact spectral triple which is $1^{+}$-summable, in particular $q$-summable for any $q>1$. More precisely, we have
$\dim\big(\mathrm{C}(\mathbb{T}),\mathrm{L}^{p}(\mathbb{T}),\tfrac{1}{\mathrm{i}}\tfrac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}\big)=1.$
Proof : As we said the operator $\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}$ generates a strongly continuous group of operators acting on $\mathrm{L}^{p}(\mathbb{T})$, namely the group of translations. By Example 2.7 and Example 2.8, we deduce that the unbounded operator $D\overset{\mathrm{def}}{=}\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}$ is bisectorial with $\omega_{\mathrm{bi}}(D)=0$ and admits a bounded $\mathrm{H}^{\infty}(\Sigma_{\omega}^{\mathrm{bi}})$ functional calculus for any angle $\omega>0$ on the Banach space $\mathrm{L}^{p}(\mathbb{T})$. By Proposition 8.3, we have $\operatorname{\mathrm{Lip}}_{D}(\mathrm{C}(\mathbb{T}))=\operatorname{\mathrm{Lip}}_{D}(\mathrm{L}^{\infty}(\mathbb{T}))\cap\mathrm{C}(\mathbb{T})=\mathrm{W}^{1,\infty}(\mathbb{T})\cap\mathrm{C}(\mathbb{T})$. Since this last space contains $\mathrm{C}^{\infty}(\mathbb{T})$, it is dense in $\mathrm{C}(\mathbb{T})$.
Finally, it is stated without proof in [Kat76, Example 6.31 p. 187] that the operator $D=\frac{1}{\mathrm{i}}\tfrac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}$ has compact resolvent. Actually, below we prove a stronger fact than the stated result in [Kat76, Example 6.31 p. 187].
For any function $f\in\mathrm{L}^{p}(\mathbb{T})$, according to [BeS22, Lemma 2 p. 24] we have
$\left\|(\mathrm{Id}+\Delta)^{-1}(f)\right\|_{\mathrm{W}^{2,p}(\mathbb{T})}\lesssim\left\|f\right\|_{\mathrm{L}^{p}(\mathbb{T})}.$
Note that $(D+\mathrm{i})^{-1}=(D-\mathrm{i})(\mathrm{Id}+\Delta)^{-1}$. If $J\colon\mathrm{W}^{1,p}(\mathbb{T})\to\mathrm{L}^{p}(\mathbb{T})$ is the canonical embedding then
(8.10) $a_{n}((D+\mathrm{i})^{-1})\leqslant a_{n}\big(J\circ(D-\mathrm{i})\circ(\mathrm{Id}+\Delta)^{-1}\big)$
@@
So, the quasi-norm can be estimated by
@@
Hence $\left\|(D+\mathrm{i})^{-1}\right\|_{p,\infty}<\infty$ for every $p\geqslant 1$. We deduce that the resolvent operator $(D+\mathrm{i})^{-1}\colon\mathrm{L}^{p}(\mathbb{T})\to\mathrm{L}^{p}(\mathbb{T})$ belongs to the space $S^{1,\infty}_{\mathrm{app}}(\mathrm{L}^{p}(\mathbb{T}))$, defined in (5.8). Thus the triple $(\mathrm{C}(\mathbb{T}),\mathrm{L}^{p}(\mathbb{T}),\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta})$ is $1^{+}$-summable.
Now, we prove that this result is optimal. When $p=2$, the statement is well known. The Banach compact spectral triple $(\mathrm{C}(\mathbb{T}),\mathrm{L}^{p}(\mathbb{T}),\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta})$ is not $1$-summable. Indeed, by [AmG16, Remark B.7 p. 127], the spectrum of the Dirac operator $D\overset{\mathrm{def}}{=}\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}$ is independent of $p$. In the case $p=2$, it is known that $\operatorname{Sp}D=\mathbb{Z}$ and that each eigenvalue is simple, see e.g., [Var10, Exercise 8.1] or [Gin09, Notes 2.1.2 p. 31]. Using Weyl’s inequality (5.12) and the spectral theorem [Haa06, Proposition A.3.1 p. 276] for the resolvent operator $R_{z}(D)\overset{\mathrm{def}}{=}(z-D)^{-1}$ , we see that for any $z\not\in\operatorname{Sp}D$ we have
$\infty=\sum_{n\in\mathbb{Z}}\frac{1}{|z-n|}=\sum_{n\in\mathbb{Z}}\lambda_{n}(R_{z}(D))=\left\|\big(\lambda_{n}(R_{z}(D))\big)\right\|_{\ell^{1}}$
@@
Consequently the resolvent operator $R_{z}(D)$ does not belong to the space $S^{1}_{\mathrm{app}}(\mathrm{L}^{p}(\mathbb{T}))$. We conclude that the triple is not $1$-summable according to Definition 7.3. Since $S^{q,\infty}_{\mathrm{app}}(\mathrm{L}^{p}(\mathbb{T}))\subset S^{1}_{\mathrm{app}}(\mathrm{L}^{p}(\mathbb{T}))$ for any $0<q<1$ the proof is complete.
@@
(8.14) $\big\langle[f],(\mathrm{L}^{p}(\mathbb{T}),\pi,F^{\prime})\big\rangle_{\mathrm{K}_{1}(\mathrm{C}(\mathbb{T})),\mathrm{K}^{1}(\mathrm{C}(\mathbb{T}),\mathscr{L}^{p})}=-\operatorname{wind}f,$
@@
Proof : First proof. The operator $|D|$ is the unbounded Fourier multiplier defined by the symbol $(|n|)_{n\in\mathbb{Z}}$ and $D$ by $(n)_{n\in\mathbb{Z}}$. The first assertion is then a consequence of (8.12). Hence, for any function $g\in\mathrm{C}^{\infty}(\mathbb{T})$, we have
@@
We now estimate the differences. For the first one, by the fundamental theorem of calculus,
@@
Second proof of the first assertion. Observe that by $|D|$ is a pseudo-differential operator of order $1$. Moreover, recall that if $f\in\mathrm{C}^{\infty}(\mathbb{T})$ then $M_{f}$ is a differential operator of order 0, hence a pseudo-differential operator of order 0 by [Dieu88, Remark 23.26.5]. Consequently, according to [RuT10, Remark 5.3.3 p. 423] the commutator $|D|,M_{f}]$ is a pseudo-differential operator of order 0. Such operator induces a bounded operator on the Banach space $\mathrm{L}^{p}(\mathbb{T})$ by [RuT10, Theorem 5.2.23 p. 420].
Remark 8.9
We can probably replace $\mathrm{C}^{\infty}(\mathbb{T})$ by the space $\mathrm{W}^{1,\infty}(\mathbb{T})$ in the previous result.
@@
Suppose that $1<p<\infty$. Note that by [EnN00, Proposition 4.10 p. 31] any multiplication operator $M_{f}\colon\mathrm{L}^{p}(\mathbb{R})\to\mathrm{L}^{p}(\mathbb{R})$, $g\mapsto fg$ by a measurable function $f\colon\mathbb{R}\to\mathbb{C}$ is bounded if and only if $f$ is essentially bounded. In this case, we have $\left\|M_{f}\right\|_{\mathrm{L}^{p}(\mathbb{R})\to\mathrm{L}^{p}(\mathbb{R})}=\left\|f\right\|_{\mathrm{L}^{\infty}(\mathbb{R})}$. Consequently, we can consider the isometric homomorphism $\pi\colon\mathrm{L}^{\infty}(\mathbb{R})\to\mathrm{B}(\mathrm{L}^{p}(\mathbb{R}))$, $f\mapsto M_{f}$. It is well known [Ren04, Corollary 14 p. 92] [Var10, Remark 7.14] that $(\mathrm{C}_{0}(\mathbb{R}),\mathrm{L}^{2}(\mathbb{R}),\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x})$ is a locally compact spectral triple, where the used homomorphism is the restriction of $\pi$ on the $\mathrm{C}^{*}$-algebra $\mathrm{C}_{0}(\mathbb{R})$. We begin to prove a an $\mathrm{L}^{p}$-variant of this classical fact. Here, we consider the closure of the closable unbounded operator $\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}\colon\mathrm{C}^{\infty}_{0}(\mathbb{R})\subset\mathrm{L}^{p}(\mathbb{R})\to\mathrm{L}^{p}(\mathbb{R})$. We denote again $\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}$ its closure whose the domain $\operatorname{dom}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}$ is the Sobolev space
$\mathrm{W}^{1,p}(\mathbb{R})\overset{\mathrm{def}}{=}\{f\in\mathrm{L}^{p}(\mathbb{R}):f\text{ is absolutely continuous},f^{\prime}\in\mathrm{L}^{p}(\mathbb{R})\},$
see, e.g., [Haa06, Proposition 8.4.1 p. 228].